40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed May 1 00:05:03 2013, Koji, Update, LSC, More stable lock of PRMI (REF33I and AS55Q) Screenshot.png
    Reply  Wed May 1 10:42:35 2013, Steve, Update, LSC, cleanup  
    Reply  Wed May 1 14:42:45 2013, Jenne, Update, LSC, POP now has lens in front of PD 
Message ID: 8518     Entry time: Wed May 1 10:42:35 2013     In reply to: 8517
Author: Steve 
Type: Update 
Category: LSC 
Subject: cleanup  

 

 Optics from the car were placed into glass door cabinet E0

ELOG V3.1.3-